Solve the system of equations using the addition (elimination) method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is "no solution" or "infinitely many solutions" and state how you arrived at that conclusion.
5x - 3y = 1/3
-10x + 6y = 2/3


Answer by ankor@dixie-net.com(22740)

5x - 3y = 1/3
-10x + 6y = 2/3
:
Multiply the 1st equation by 6, and the 2nd equation by 3, this gets rid of the
denominator and give the coefficient of y the same in both equations
+30x - 18y = 2
-30x + 18y = 2
:
you can see that this has "no solution"